OpenStudy (anonymous):

What is the area of a circle with radius 7.5?

OpenStudy (anonymous):

\[\pi r^2\]

OpenStudy (anonymous):

\[625\pi \div4\]

OpenStudy (anonymous):

pi*r^2 7.5^2=56.25 pi*56.25=176.7144

OpenStudy (anonymous):

area = πr^2 = 3.14 * (7.5)^2 = 3.14*56.25 = 176.63 if u use π = 22/7, then area = πr^2 = (22/7)*(7.5)^2 = (22/7)*56.25 = 176.79